OpenStudy (anonymous):

One leg of a right triangle measures 8 inches and the hypotenuse 17 inches. Determine the length of the other leg. a. 9 inches b. 13 inches c. 15 inches d. 17 inches

OpenStudy (rea201):

You are going to be usisng the pythagorean theorem

OpenStudy (rea201):

You can't give straight answers it is against the coc

OpenStudy (rea201):

Pythagorean theorem is a^2+b^2=c^2 where a and b are the length of the legs and c is the length of the hypotnuese

OpenStudy (rea201):

so you have 8^2+b^2=17^2

OpenStudy (rea201):

so next you get b=sqrt(17^2-8^2)

OpenStudy (anonymous):

Its 9 because the sum of the lengths of the two legs can't be greater than the hypotenuse so you would just do 17-8 and you would get 9.I think that's right but i am not 100% sure.

OpenStudy (anonymous):

I think it's 15..

OpenStudy (rea201):

drsuz that is incorrect you use the pythagorean theorem as i showed above. If you had the full perimeter than you could subtract 17 and 8 of the that but other wise the only way to do it with the information given is the pythagorean theorem

OpenStudy (rea201):

It is 15

OpenStudy (anonymous):

Yeah I kinda figured..I realized that I did the complete wrong math..lol..sorry @Rea201

OpenStudy (rea201):

That is fine. :)

OpenStudy (anonymous):

Yay!!!